You are a business owner of a firm that services trucks. A customer would like to rent a truck from you for one week, while you service his truck. You must decide whether or not to rent him a truck. You have an extra truck that you will not use for any other purpose during this week. This truck is leased for a full year from another company for $350/ week plus $.50 for every mile driven. You also have paid an annual insurance premium, which costs $70/ week to insure the truck. The truck has a full 100-gallon fuel tank. The customer has offered you $500 to rent the truck for a week. The price includes the 100 gallons of fuel that is in the tank. It also includes up to 500 miles of driving. The customer will pay $.50 for each additional mile that he drives above the 500 miles. You anticipate that the customer will bring back the truck with an empty fuel tank and will have driven more than 500 miles. You sell fuel to truckers at a retail price $3.35/gallon. Any fuel you sell or use can be replaced at a wholesale price of $2.95/gallon. The customer will rent a truck from another company if you do not accept the proposed deal. In either case, you will service his truck. You know the customer and are confident that he will pay all charges incurred under the agreement. 1. Should you accept or reject the proposed deal? Why, or why not? Show calculations. 2. Would your answer change if your fuel supplier limited the amount of fuel up to 100 gallons/ week you could purchase from him at the wholesale price? Explain.
Answer #1

Fixed cost

Insurance cost = $70/week

Leasing cost = $350/week

Fixed cost is not to be used for decision making purpose as even if the truck is not rented, fixed cost would have to be incurred

Variable cost

500 miles of driving costs = 500*0.5 = $250. Any additional mile of drive would not make any difference as customer will pay $0.5 which is the same amount I need to pay to the company from where the truck has been leased.

Profit by selling 1 gallon fuel = ($3.35-$2.95) = $.4

Opportunity cost of renting and giving away profit of 100 gallon which is 100*($3.35-$2.95) = $40

Cost of 100 gallon fuel = 100*2.95 = $295

So total variable cost of renting out the truck = Driving cost + profit of selling 100 gallon fuel +cost of 100 gallon fuel = $250+$40+295 = $585

  1. Truck Should not be rented as variable cost ($585) is more than what he is offering ($500)
  2. If fuel supplier limits fuel up to 100gallons/week, that would not change the decision as that would not lower my variable cost